Tohban Report 2012-11-27
From RHESSI Wiki
|Start Date:||21 Nov 2012|
|End Date:||27 Nov 2012|
|Next Tohban:||Hugh Hudson|
|List all reports|
After the M flares of last week, solar activity dropped about a week ago (coincident with this tohban shift!) and is now slowly picking back up. Currently there are 4 active regions on the disk, one of which (AR 11620) is producing several C flares and 1 M flare today. Activity is likely to continue at this level, as STEREO B does not show any large, complex ARs about to rotate onto the disk.
How many GOES flares occurred?
Flares above B, C, M, X class were 8 29 2 0
And how many of these are listed in the RHESSI flare list?
Flares above B, C, M, X class were 5 18 2 0
And how many had EXCELLENT coverage?
Flares above B, C, M, X class were 2 7 0 0
Memory management is ok, with the SSR reaching a max of 22% and emptying at the end of each pass set. Decimation remains at normal/vigorous, and we are not collecting events at night.
There were no significant changes in the cold plate temperatures or cryocooler power.
There was one data gap (20 mins around 2012-11-22T05:00:00.000), with the data later filled in but showing some bad packets.
Detectors: there was one isolated spike in D4, which also gets a little spiky during the SAA, but nowhere near the wild behavior of D8. Other than that there was no unusual detector activity -- D8 is still showing increased counts coming out of the SAA and D2 is showing a light leak problem. Both of these are known issues and do not need correction at this time.
There were no changes to the detector settings this week.
|HLAT Decimation||Rear decimation weight 6, no front decimation|
|Night time data (fronts)||+/- 4 minutes|
|Night time data (rears)||+/- 4 minutes|
|Require extra passes?||No|
|Requirement for moving pointer?||No|